OpenStudy (anonymous):

Solve the following system of equations: x − 2y = 14 x + 3y = 9 A (1, 12) B (–1, −12) C (12, –1) D (12, 1)

OpenStudy (anonymous):

Put the value of x from 2nd equation in 1st equation 2(2y - 1) + y = 3 4y - 2 + y = 3 5y = 3 + 2 5y = 5 y = 5/5 y = 1 x = 2y - 1 x = 2(1) - 1 x = 2 -1 x = 1

OpenStudy (whpalmer4):

Either solve one of the equations for x in terms of y, then plug that into the other equation (substitution), or subtract one equation from the other, giving you an equation only in terms of one of the variables (elimination)

OpenStudy (anonymous):

hence, x = 1 and y = 1

OpenStudy (aravindg):

@ObeyMfnBino PLEASE donot provide direct solution like that . It is against openstudy COC

OpenStudy (aravindg):

thank you

OpenStudy (anonymous):

x − 2y = 14 ---> times 3 x + 3y = 9 -----> times 2 3x-6y=42 2x+6y=18 --------- 5x =60 x=12 Do you see how I got that? Now, plug the value of x into one of the original equations. :o)

OpenStudy (anonymous):

people do it to me all the time.. but i respect your terms wont ever agian, Your Welcome

OpenStudy (whpalmer4):

\[x-2y = 14\]\[x - 2y - (x + 3y) = 14 - 9\] (this is subtracting one equation from the other)\[x - 2y -x - 3y = 5\]\[-5y = 5\]\[y = -1\]\[x - 2(-1) = 14\]\[x + 2 = 14\]\[x = 12\] Check work: \[12 - 2(-1) = 14\]\[12 + 3(-1)= 9\]

OpenStudy (whpalmer4):

[The solution already having been revealed, I don't see any harm in illustrating a different approach]

OpenStudy (anonymous):

So the answer would be C, right?

OpenStudy (whpalmer4):

As an old math teacher was fond of saying, "is that an answer, or a prayer?" :-) Look at the answers you have to choose from. Try them out in the system of equations. Be confident in your choice.

OpenStudy (anonymous):

I'm going with C